OpenStudy (anonymous):

Is (3, 13) a solution of the equation y = 4x? (1 point) Explain? :/

jimthompson5910 (jim_thompson5910):

any point is (x,y) so (3, 13) means x = 3 and y = 13

jimthompson5910 (jim_thompson5910):

from here, plug in x = 3 and y = 13 y = 4x 13 = 4*3 Keep going...

OpenStudy (anonymous):

How do you know that y=13 and x = 3? Sorry i'm bad at math

jimthompson5910 (jim_thompson5910):

because the point is (3,13)

jimthompson5910 (jim_thompson5910):

(x, y) = (3,13) so x = 3 y = 13

OpenStudy (anonymous):

its y,x though.?

OpenStudy (anonymous):

y = 4x

jimthompson5910 (jim_thompson5910):

no, any point is of the form (x,y)

jimthompson5910 (jim_thompson5910):

x comes first

OpenStudy (anonymous):

Oh Okay, What Do We Do After That?

jimthompson5910 (jim_thompson5910):

y = 4x 13 = 4*3 What's the next step

OpenStudy (anonymous):

Uhm.. well it will end up being 13 = 12 so divide?

jimthompson5910 (jim_thompson5910):

good does 13 equal 12 ?

OpenStudy (anonymous):

Its No.. Right!? :D

jimthompson5910 (jim_thompson5910):

that's correct, since 13 = 12 is FALSE, this means that (3,13) is NOT a solution

jimthompson5910 (jim_thompson5910):

it's only a solution if x and y make the equation true

OpenStudy (anonymous):

So in order for it to be true it will have to be 13 = 13?

jimthompson5910 (jim_thompson5910):

yes or SOME NUMBER = SAME NUMBER

OpenStudy (anonymous):

it feels so good to finally understand one equation of algebra, Do you have skype? aha

jimthompson5910 (jim_thompson5910):

ex: (3,12) is a solution y = 4x 12 = 4*3 12 = 12 and so is (8,32) y = 4x 32 = 4*8 32 = 32 etc etc There are infinitely many of these ordered pairs

OpenStudy (anonymous):

I got you

jimthompson5910 (jim_thompson5910):

that's great

OpenStudy (anonymous):

can you keep helping me? c; lol

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

Which ordered pair is a solution of the equation y = 3x? (1 point) So would (y) = 1 since its not given a number?

OpenStudy (anonymous):

here are the answers (–2, –9) (–8, –18) (–8, –3) (–10, –30)

jimthompson5910 (jim_thompson5910):

Do the same thing, but now we're talking about a different equation y = 3x -9 = 3(-2) .... Plug in x = -2 , y = -9 -9 = -6 which is FALSE...so (-2, -9) is NOT a solution

OpenStudy (anonymous):

This is hard. I'm confused.

OpenStudy (anonymous):

So you keep going through the answers to see which one is the same?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

you need to test them all

OpenStudy (anonymous):

Plug in x = -2 , y = -9 -9 = -6 How did you get -6?

jimthompson5910 (jim_thompson5910):

3 times -2 is -6

jimthompson5910 (jim_thompson5910):

y = 3x -9 = 3 times -2 -9 = -6

OpenStudy (anonymous):

Oh I Got You!

jimthompson5910 (jim_thompson5910):

ok great